# Disruptions of stars and binary systems on chaotic orbits in the axisymmetric Milky Way center

### v1

@@ 6.1 Relaxation times
Perhaps a larger concern is that this is a local measure, at some position $\mathbf{r}$, while we would ideally like to compare to the value for a given orbit (characterised by energy $\epsilon$) whose trajectory may take it through markedly different local conditions over the timescale in which relaxation acts. However, scattering is generally most effective at apoapse (where the orbit spends more time and velocity distributions are lower) [^12] This is not always true, for example very steep density profiles will may have enough stars at small radii to make periapse dominate the scattering process (Stone18). However the assumption holds well for most observed stellar systems and it is possible that systems for which it is not true will evolve until it is true (as effective scattering at small radii will lead to redistribution of energy outwards, moving stars to larger radii and rebalancing the central excess)..
Our orbits do not have a single apoapse, but their typical apoapse is of order the (cylindrical) circular radius associated with their energy. Thus we calculate the number densities at $R=R_{\rm c}(\epsilon)$ (and $z=0$, ignoring the slight decrease in density away from the disk plane) and similarly we approximate the velocity dispersion as the circular velocity at that radius, i.e.
@@
We also show an alternate evaluation of the energy relaxation time including a simple treatment of the presence of massive perturbers in the outer GC (for example giant molecular clouds of masses > $10^{5}M_{\odot}$ which might significantly increase the rate of energy redistribution). This effect is detailed fully in section 6.5.1.
@@ 6.2 Full and empty diving regimes
In the empty loss wedge case, typical disrupted systems will have $|\gamma_{z}|$ close to one, as they would not have had time to diffuse to lower values.
@@
It is also possible for angular momentum redistribution to move stars in $l_{\theta}$, thus allowing a higher fraction of orbits than $F_{\rm dive}$ to disrupt, as previously stable orbits may be scattered onto chaotic trajectories. We do not include this effect in our models, but note that if it is maximally efficient it will lead to $F_{\rm dive}\rightarrow F_{\rm phys}$ and thus increase the rates by a factor of a few.
@@ 6.3 Disruption rates from diving orbits
The final ingredient needed to calculate the collisionless disruption rate from diving orbits is the number of systems that we can expect to have a given energy and angular momentum. We calculate this by numerically sampling the distribution functions, as detailed in appendix A. Now we can combine the approximated number densities in table 1 with equations 53 and 64 to write down the relevant form of equation 58:
@@
where $\bar{N}_{\rm I}$ is the total number of systems in component I. $N_{\rm I}=M_{\rm I}/\langle m_{\rm s}\rangle$ where $M_{\rm I}$ is the total mass of the component, and we have already taken the average mass of test particles to be $\langle m_{\rm s}\rangle=1M_{\odot}$. The only components that we consider here are the NSC and NSD, and we will see that even the latter contributes little to the rates, thus justifying why we ignore the contribution of the bulge.
In Figure 11, we show the stellar concentrations [^13] We note that the dimensionless function $n(\rho,\gamma_{z})$ gives the number density of stars in the dimensionless space of ($\rho,\gamma_{z}$). To avoid confusion with the true (dimensional) 3D number density of stars $n_{\rm s}$ (as is e.g. plotted in figure 10), we term $n(\rho,\gamma_{z})$ to be the stellar “concentration.” $n(\rho,\gamma_{z})$, timescales $t_{\rm max}$ and relative disruption rates $\bar{\Gamma}_{\rm d}$ for five values of $l_{\rm t}$ (assuming $t_{\rm gal}$ is large). These can be mapped to approximate values of the tidal disruption radius $r_{\rm t}$ through equation 23, and range from $\approx$0.5 AU to $\approx$5000 AU (in multiples of $\sim$10). Thus for a $\sim 10^{6}M_{\odot}$ BH, the smallest $l_{\rm t}$ shown is appropriate for the tidal disruption of a Sun-like star, the middle value for a relatively tight ($\sim$ 0.5 AU) binary or a giant star, and the largest for a wide ($\sim$ 50 AU) binary or a typical triple star system.
@@
For most of the parameter space shown, the longer (and thus rate-limiting) timescale is $t_{\rm D}$, the time taken for a single diving orbit to come sufficiently close to disrupt. We show the line for which $t_{\rm D}=t_{\mathrm{rel},l_{z}}$ in black, highlighting that we are in the full loss wedge regime for all but smallest radii and largest $|\gamma_{z}|$. We can see why the empty regime dominates at small radii by expressing the relaxation time in terms of $T_{\rm c}$ ($t_{\mathrm{rel},\epsilon}\propto T_{\rm c}^{-3}$ while $t_{\rm D}\propto T_{\rm c}^{-1}$) and remembering that $T_{\rm c}$ increases rapidly with radius. We know that $t_{\mathrm{rel},l_{z}}\propto\gamma_{z}^{2}\rightarrow 0$ as $|\gamma_{z}|\rightarrow 0$. Meanwhile, $t_{\rm D}\propto\frac{1}{1-|\gamma_{z}|}$ gets very large as $|\gamma_{z}|\rightarrow 1$. This explains the shape of the transition line in these two extremes, and thus why the empty regime does not extend to low $|\gamma_{z}|$.
@@
Generally, the peaks in the differential rate $\frac{\partial^{2}\bar{\Gamma}_{\rm d}}{\partial\rho\partial\gamma_{z}}$ correspond to combinations of high $n(\rho,\gamma_{z})$ and low $t_{\rm max}$. There is an effective maximum $\rho$ ($\sim$ 1) above which $t_{\rm max}$ is prohibitively large, and the differential rate falls off quickly. The total rate, which can be seen in the $\bar{\Gamma}_{\rm d}$ panels of Figure 11, is generally dominated by systems in the full regime, and prefers systems from moderate ($\rho\sim 0.5$) radii for which $n$ is significant. For most values of $l_{\rm t}$, the total disruption rate is overwhelmingly dominated by the full loss wedge regime.
@@
Almost all disrupted systems have low $|\gamma_{z}|$ (due to the shorter $t_{\rm max}$ values for low $l_{z}$). While a wide range of disruption parameters, such as the inclination and orbital eccentricity (see equations 51 and 52 respectively) are possible, the most likely scenario will be a near-parabolic, grazing disruption strongly misaligned with the disk plane [^14] Both photometric (Nishiyama & Schödel, 2013) and dynamical (Sormani et al., 2022) studies show that the NSD, and hence the whole GC, is well aligned with the Milky Way stellar disk.. We therefore expect to see more ejecta (tidal debris for TDEs; hypervelocity stars for the Hills mechanism) directed along the polar directions orthogonal to the galactic plane.
@@
This notable injection anisotropy would have been reversed had we found that most disruptions occur in the empty loss wedge regime; in this case, almost all disruptions would occur with $|\gamma_{z}|$ close to 1, and $l_{\theta,\rm p}=0$, limiting ejecta to fly out within the equatorial plane of the galaxy.
@@
Figure 11: Dimensionless stellar concentrations (top row), rate-limiting timescales (second row), differential contribution to relative rates (third row) and cumulative relative rates (fourth row) for collisionless disruptions by diving orbits for a GC model composed entirely of systems of a given $l_{\rm t}$. Different $l_{\rm t}$ values are labelled in units of [pc2 Myr-1], and are shown in different columns. All panels plot their respective quantities against $\rho=\log_{10}(R_{\rm c}(\epsilon)/{\rm pc})$ and $\gamma_{z}=l_{z}/l_{\rm t}$. Note that the colour bar legends vary from panel to panel. The solid black lines show the transition from empty to full loss region (with the full regime, $t_{\mathrm{rel},l_{z}}<t_{\rm d}$, predominating to the right of the black lines at large $\rho$). Though the stellar concentrations show very little variation with $|\gamma_{z}|$, the total number with $|\gamma_{z}|\leq 1$ is proportional to $\lambda_{\rm t}$, and this increase leads systems with larger $l_{\rm t}$ to have a significantly larger rate. The final row shows the cumulative rate for all values of $\rho$ and $\gamma_{z}$ less than or equal to that at each point, thus the total rate is the value at the top right corner. From the final row, we see that disruptions come overwhelmingly from the full loss wedge regime.
@@ 6.4 Disruption rates from scattering
Systems can also acquire low angular momentum orbits by the collisional mechanism of scattering alone (i.e. ignoring any collisionless evolution in $l$ produced by asphericity of the potential). This rate now depends on a much smaller region of phase space: $l<l_{\rm t}$ instead of $l_{z}<l_{\rm t}$ (i.e. a small volume in 3 dimensions, instead of just 1).Collisional scattering of stars into the loss cone of a spherical star system has been widely studied in the TDE literature (Frank & Rees, 1976; Cohn & Kulsrud, 1978; Magorrian & Tremaine, 1999; Wang & Merritt, 2004; Stone & Metzger, 2016) and thus for comparison we’ll present a brief and simple version of the calculation, mostly drawn from Merritt (2013a) chapter 6. In this calculation there are analogous full and empty regimes, now dependent on whether the diffusive change in angular momentum over one period is more (full) or less (empty) than $l_{\rm t}$.
@@ 6.4.1 Full loss cone
We begin with the full loss cone case where radial periods, $T_{r}$, are much longer than the timescale over which angular momentum is redistributed, $t_{\mathrm{rel},l}$. The particular angular momentum relaxation time we care about is that at the loss cone boundary, where $l=l_{\rm t}$. The full loss cone regime is attained at larger radii, where the loss cone shrinks to a small “pinhole” in angular momentum space. In this regime, neglible depletion occurs, and the relevant version of equation 58 (for a given $l_{\rm t}$) is
@@
where $\bar{N}_{\rm I}$ is the total number of systems in component I (the above assumes $b>-1$, which is true for our fitted models). To compare directly with the empty loss cone rate we can integrate equation 67 over angular momenta to find
@@ 6.4.2 Empty loss cone
In the empty loss cone regime, systems are disrupted faster than they are repopulated (i.e. $T_{r}\ll t_{\mathrm{rel},l}$). We therefore need to account for the fact that the relevant stellar concentration is reduced, i.e. $n_{E}(\rho,\gamma)\leq n(\rho,\gamma)$. The classic result in loss cone theory is to find a stationary solution to the Fokker-Planck equation in angular momentum space (Cohn & Kulsrud, 1978). In spherical symmetry, this solution features logarithmic depletion of the distribution function at smaller values of $l$. In the limit of $T_{r}\ll t_{\rm rel,l}$, the distribution function goes to zero at $l=l_{\rm t}$. In other words, stars random walk through angular momentum space, but do so with such small steps that they are destroyed on the first orbit for which $l\leq l_{\rm t}$.
Once the stationary solution in the empty loss cone regime is established, the disruption rate becomes
@@
One way of interpreting this result is to say that for a fixed energy, a fraction $1/|\ln(\lambda_{\rm t}^{2})|$ of all systems pass into the loss cone over each energy relaxation time $t_{\mathrm{rel},\epsilon}$. For practical evaluation of this formula, we find $n(\epsilon)$ by sampling from the DF in AGAMA and summing over components. This procedure is detailed in appendix A.
At the transition point between the full and empty loss cone regimes (where $t_{\mathrm{rel},\epsilon}\sim T_{\rm c}$) the rate should smoothly interpolate between the full and empty limits, but for simplicity here we will use
@@
In general, the empty loss cone regime dominates at small radii (where orbital periods are small) and the full loss cone regime comes into effect at larger radii.
@@ 6.5 Comparing rates
In Figure 12, we now compare the total contribution from scattering to that from diving orbits.
Most scattered systems are in the empty loss-cone regime (the exception being for the lowest $l_{\rm t}$, i.e. TDEs). This result may be partially spurious due to our neglect of bulge stars. This artificially reduces the star-star relaxation rate at large radii ($\rho\gtrsim 2$), making the dominance of the empty loss cone here likely artificial; however we can also see that even a full loss cone at such large radii would make little difference to the total rates.
We also find that we are in the full loss cone regime for low $l_{\rm t}$ around $r_{\rm infl}$, and furthermore predict a substantially lower overall rate for scattering-driven TDEs, $10^{-5}$ yr-1, than commonly reported values of order $10^{-4}$ yr-1 (see for example Generozov et al. 2018). This difference comes from explicitly modelling the anisotropic distribution of $l$, as detailed by equation 69. For the NSC we find a steeper dependence on $l$ than the linear law appropriate for isotropic systems (see Figure 20), which suppresses the number of low $l$ systems and thus the overall rate. In other words, our NSC model has a significant tangential bias at small radii, which has been explored in other contexts as a way to reduce the full loss cone TDE rate (Lezhnin & Vasiliev, 2015; Teboul et al., 2024). This result evidently depends on the NSC DF used, and careful observations or modelling of the distribution of low $l$ NSC orbits would be needed to confirm this suppression.
@@
Figure 12: The contribution to the relative rate of disruption (top panels) and cumulative relative rates (bottom panels) for disruptions by diving orbits (left panels) and scatterings (right panels) - for systems of a given $l_{\rm t}$ (in units of [pc2 Myr-1]). We also give the approximate tidal radius ($r_{\rm t}\sim l_{\rm t}^{2}(2GM_{\rm BH})^{-1}$ where $M_{\rm BH}=4\cdot 10^{6}M_{\odot}$). For the scattering calculations we show the empty loss cone rate (solid line) and the full loss cone rate (dotted). We take the true rate of disruptions by scattering to be whichever is smaller, and in the cumulative rate the solid or dotted regions correspond to which rate is relevant in that region. Vertical lines show the $\rho$ which contributes half of the relative rate (giving a rough idea of where we expect disrupted systems to originate from).

### v3

@@ 6.1 Relaxation times
Perhaps a larger concern is that this is a local measure, at some position $\mathbf{r}$, while we would ideally like to compare to the value for a given trajectory (characterised by energy $\epsilon$) whose orbit may take it through markedly different local conditions over the timescale in which relaxation acts. However, scattering is generally most effective at apoapse (where the orbit spends more time and velocity distributions are lower). [^12] This is not always true: for example, very steep density profiles will may have enough stars at small radii to make periapse dominate the scattering process (Stone et al., 2018). However the assumption holds well for most observed stellar systems and it is possible that systems for which it is not true will evolve until it is true (as effective scattering at small radii will lead to redistribution of energy outwards, moving stars to larger radii and rebalancing the central excess).
Our trajectories do not have a single apoapse, but their typical apoapse is of order the (cylindrical) circular radius associated with their energy. Thus we calculate the number densities at $R=R_{\rm c}(\epsilon)$ (and $z=0$, ignoring the slight decrease in density away from the disk plane) and similarly we approximate the velocity dispersion as the circular velocity at that radius, i.e.
@@
We also show an alternate evaluation of the energy relaxation time including a simple treatment of the presence of massive perturbers in the outer GC (for example, giant molecular clouds of masses > $10^{5}M_{\odot}$ which might significantly increase the rate of energy redistribution). This effect is detailed fully in section 6.5.1.
@@ 6.2 Full and empty diving regimes
In the empty loss-wedge case, typical disrupted systems will have $|\gamma_{z}|$ close to one, as they would not have had time to diffuse to lower values. A fuller calculation would include the depletion of low $l_{z}$ systems in $n(\epsilon,l_{z})$, which as shown in Magorrian & Tremaine (1999) reduces the empty loss-wedge rate by a factor of order $F_{\rm dive}$. However, as we will go on to show, the total number of disruptions is dominated by the full loss wedge and thus this correction should only have a small effect.
@@
It is also possible for angular momentum redistribution to move stars in $l_{\theta}$, thus allowing a higher fraction of orbits than $F_{\rm dive}$ to disrupt, as previously stable orbits may be scattered onto chaotic trajectories. We do not include this effect in our models, but note that if it is maximally efficient it will lead to $F_{\rm dive}\rightarrow F_{\rm phys}$ and thus increase the rates by at most a factor of a few.
@@ 6.3 Disruption rates from diving orbits
The final ingredient needed to calculate the collisionless disruption rate from diving orbits is the number of systems that we can expect to have a given energy and angular momentum. We calculate this by numerically sampling the distribution functions, as detailed in appendix A. Now we can combine the approximated number densities in table 1 with equations 53 and 65 to write down the relevant form of equation 59:
@@
where $\bar{N}_{\rm I}$ is the total number of systems in component I. $N_{\rm I}=M_{\rm I}/\langle m_{\rm s}\rangle$ where $M_{\rm I}$ is the total mass of the component, and we have already taken the average mass of test particles to be $\langle m_{\rm s}\rangle=1M_{\odot}$. The only components that we consider here are the NSC and NSD, and we will see that even the latter contributes little to the rates (further justifying ignoring the contribution of systems in the bulge).
In Figure 11, we show the stellar concentrations [^13] We note that the dimensionless function $n(\rho,\gamma_{z})$ gives the number density of stars in the dimensionless space of ($\rho,\gamma_{z}$). To avoid confusion with the true (dimensional) 3D number density of stars $n_{\rm s}$ (as is e.g. plotted in figure 10), we term $n(\rho,\gamma_{z})$ to be the stellar “concentration.” $n(\rho,\gamma_{z})$, timescales $t_{\rm max}$ and relative disruption rates $\bar{\Gamma}_{\rm d}$ for five values of $l_{\rm t}$ (using $t_{\rm gal}=10$ Gyr). These can be mapped to approximate values of the tidal disruption radius $r_{\rm t}$ through equation 23, and range from $\approx$0.5 AU to $\approx$5000 AU (in multiples of $\sim$10). Thus for a $\sim 10^{6}M_{\odot}$ BH, the smallest $l_{\rm t}$ shown is appropriate for the tidal disruption of a Sun-like star, the middle value for a relatively tight ($\sim$ 0.5 AU) binary or a giant star, and the largest for a wide ($\sim$ 50 AU) binary or a typical triple star system.
@@
For most of the parameter space shown, the longer (and thus rate-limiting) timescale is $t_{\rm D}$, the time taken for a single diving orbit to come sufficiently close to disrupt. We show the line for which $t_{\rm D}=t_{\mathrm{rel},l_{z}}$ in black, highlighting that we are in the full loss-wedge regime for all but smallest radii and largest $|\gamma_{z}|$. We can see why the empty regime dominates at small radii by comparing the relatively shallow increase in the relaxation time with radius (see figure 10) while $t_{\rm D}\propto T_{\rm c}$ increases rapidly in comparison. We know that $t_{\mathrm{rel},l_{z}}\propto\gamma_{z}^{2}\rightarrow 0$ as $|\gamma_{z}|\rightarrow 0$. Meanwhile, $t_{\rm D}\propto\frac{1}{1-|\gamma_{z}|}$ gets very large as $|\gamma_{z}|\rightarrow 1$. This explains the shape of the transition line in these two extremes, and thus why the empty regime does not extend to low $|\gamma_{z}|$.
@@
Generally, the peaks in the differential rate $\frac{\partial^{2}\bar{\Gamma}_{\rm d}}{\partial\rho\partial\gamma_{z}}$ correspond to combinations of high $n(\rho,\gamma_{z})$ and low $t_{\rm max}$. There is an effective maximum $\rho$ ($\sim$ 1) above which $t_{\rm max}$ is prohibitively large, and the differential rate falls off quickly. The total rate, which can be seen in the $\bar{\Gamma}_{\rm d}$ panels of Figure 11, is generally dominated by systems in the full regime, and prefers systems from moderate ($\rho\sim 0.5$) radii for which $n$ is significant. For most values of $l_{\rm t}$, the total disruption rate is overwhelmingly dominated by the full loss-wedge regime.
@@
Almost all disrupted systems have low $|\gamma_{z}|$ (due to the shorter $t_{\rm max}$ values for low $l_{z}$). While a wide range of disruption parameters, such as the inclination and orbital eccentricity (see equations 51 and 52 respectively) are possible, the most likely scenario will be a near-parabolic, grazing disruption strongly misaligned with the disk plane. [^14] Both photometric (Nishiyama & Schödel, 2013) and dynamical (Sormani et al., 2022) studies show that the NSD, and hence the whole GC, is well aligned with the Milky Way stellar disk. We therefore expect to see more ejecta (tidal debris for TDEs; hypervelocity stars for the Hills mechanism) directed along the polar directions orthogonal to the galactic plane.
@@
This notable injection anisotropy would have been reversed had we found that most disruptions occur in the empty loss-wedge regime; in this case, almost all disruptions would occur with $|\gamma_{z}|$ close to 1, and $l_{\theta,\rm p}=0$, limiting ejecta to fly out within the equatorial plane of the galaxy.
@@
Figure 11: Dimensionless stellar concentrations (top row), rate-limiting timescales (second row), differential contribution to relative rates (third row) and cumulative relative rates (fourth row) for collisionless disruptions by diving orbits for a GC model composed entirely of systems of a given $l_{\rm t}$. Different $l_{\rm t}$ values are labelled in units of [pc2 Myr-1], and are shown in different columns. All panels plot their respective quantities against $\rho=\log_{10}(R_{\rm c}(\epsilon)/{\rm pc})$ and $\gamma_{z}=l_{z}/l_{\rm t}$. Note that the colour bar legends vary from panel to panel. The solid black lines show the transition from empty to full loss region (with the full regime, $t_{\mathrm{rel},l_{z}}<t_{\rm d}$, predominating to the right of the black lines at large $\rho$). Though the stellar concentrations show very little variation with $|\gamma_{z}|$, the total number with $|\gamma_{z}|\leq 1$ is proportional to $\lambda_{\rm t}$, and this increase leads systems with larger $l_{\rm t}$ to have a significantly larger rate. The final row shows the cumulative rate for all values of $\rho$ and $\gamma_{z}$ less than or equal to that at each point, thus the total rate is the value at the top right corner. From the final row, we see that disruptions come overwhelmingly from the full loss-wedge regime.
@@ 6.4 Disruption rates from scattering
Systems can also acquire low angular momentum orbits by the collisional mechanism of scattering alone (i.e. ignoring any collisionless evolution in $l$ produced by asphericity of the potential). This rate now depends on a much smaller region of phase space: $l<l_{\rm t}$ instead of $l_{z}<l_{\rm t}$ (i.e. a small volume in 3 dimensions, instead of just 1). Collisional scattering of stars into the loss cone of a spherical star system has been widely studied in the TDE literature (Frank & Rees, 1976; Cohn & Kulsrud, 1978; Magorrian & Tremaine, 1999; Wang & Merritt, 2004; Stone & Metzger, 2016) and thus for comparison we’ll present a brief and simple version of the calculation, mostly drawn from Merritt (2013a) chapter 6. In this calculation there are analogous full- and empty-loss cone regimes, now dependent on whether the diffusive change in angular momentum over one period is more (full) or less (empty) than $l_{\rm t}$.
@@ 6.4.1 Full loss cone
We begin with the full loss-cone case where radial periods, $T_{r}$, are much longer than the timescale over which angular momentum is redistributed, $t_{\mathrm{rel},l}$. The particular angular momentum relaxation time we care about is that at the loss-cone boundary, where $l=l_{\rm t}$. The full loss-cone regime is attained at larger radii, where the loss cone shrinks to a small “pinhole” in angular momentum space. In this regime, neglible depletion occurs, and the relevant version of equation 59 (for a given $l_{\rm t}$) is
@@
where $\bar{N}_{\rm I}$ is the total number of systems in component I (the above assumes $b>-1$, which is true for our fitted models). To compare directly with the empty loss-cone rate we can integrate equation 68 over angular momenta to find
@@ 6.4.2 Empty loss cone
In the empty loss-cone regime, systems are disrupted faster than they are repopulated (i.e. $T_{r}\ll t_{\mathrm{rel},l}$). We therefore need to account for the fact that the relevant stellar concentration is reduced, i.e. $n_{E}(\rho,\gamma)\leq n(\rho,\gamma)$. The classic result in loss-cone theory is to find a stationary solution to the Fokker-Planck equation in angular momentum space (Cohn & Kulsrud, 1978). In spherical symmetry, this solution features logarithmic depletion of the distribution function at smaller values of $l$. In the limit of $T_{r}\ll t_{\rm rel,l}$, the distribution function goes to zero at $l=l_{\rm t}$. In other words, stars random walk through angular momentum space, but do so with such small steps that they are destroyed on the first orbit for which $l\leq l_{\rm t}$.
Once the stationary solution in the empty loss-cone regime is established, the disruption rate becomes
@@
One way of interpreting this result is to say that for a fixed energy, a fraction $1/|\ln(\lambda_{\rm t}^{2})|$ of all systems pass into the loss cone over each energy relaxation time $t_{\mathrm{rel},\epsilon}$. For practical evaluation of this formula, we find $n(\rho)$ by sampling from the DF in AGAMA and summing over components. This procedure is detailed in appendix A.
At the transition point between the full and empty loss-cone regimes (where $t_{\mathrm{rel},\epsilon}\sim T_{\rm c}$) the rate should smoothly interpolate between the full and empty limits, but for simplicity here we will use
@@
In general, the empty loss-cone regime dominates at small radii (where orbital periods are small) and the full loss-cone regime comes into effect at larger radii.
@@ 6.5 Comparing rates
In Figure 12, we now compare the total contribution from scattering to that from diving orbits. We see that diving orbits have a higher relative rate of disruptions, marginally so for low $l_{\rm t}$ but boosted by more than an order of magnitude for $l_{\rm t}\gtrsim 1$.
To give an intuitive explanation for this we can evaluate both rates under the assumption of isotropy: $n(\rho,l)\propto ln(\rho)$ and $n(\rho,l_{z})\propto n(\rho)$. The number of stars in the loss cone (integrating up to $l_{\rm t}$) is then $n_{\rm LC,ISO}\sim\lambda_{t}^{2}n(\rho)$ and similarly for the loss wedge $n_{\rm LW,ISO}\sim\lambda_{\rm t}n(\rho)$. The full loss-cone rate under this assumption is approximately
$\frac{\partial\bar{\Gamma}_{\rm LC,ISO}}{\partial\rho}\sim\frac{n_{\rm FLC,ISO}}{T_{\rm c}}=\frac{\lambda^{2}_{\rm t}}{T_{\rm c}}n(\rho).$ (74)
and similarly the full loss-wedge rate is
$\frac{\partial\bar{\Gamma}_{\rm FLW,ISO}}{\partial\rho}\sim F_{\rm dive}\frac{n_{\rm LW,ISO}}{t_{D}}=\frac{\lambda^{2}_{\rm t}}{T_{\rm c}F_{\rm phys}}n(\rho),$ (75)
i.e. the two are essentially the same (as $F_{\rm phys}$ is generally close to 1). The full regime (in both cases) represents the maximum possible rate, and can only be maintained if the rate of refilling matches or exceeds the rate of depletion.
One difference between the two mechanisms of disruption comes from the fact that $t_{\rm D}$ is long compared to $t_{\mathrm{rel},l_{z}}$ down to small radii, and thus the full loss wedge extends to more systems (lower $\rho$) than the full loss cone. Another way to state this is that both mechanisms are able to deplete systems at the same rate, but the loss wedge represents a larger volume and surface are of parameter space and thus is refilled quicker.
The second difference comes from deviations from isotropy. This is most clearly seen in the power law dependence of $n(\rho,l)$ on $l$, with a tangential bias (fewer radial orbits) leading to a steeper power law and thus potentially $n_{\rm LC}\ll n_{\rm LC,ISO}$. A tangential bias is seen in our GC model (see section 6.4.1) and confirmed in recent observations of the GC (Feldmeier-Krause et al., 2025) and has been explored in other contexts as a way to reduce the full loss-cone TDE rate (Lezhnin & Vasiliev, 2015; Teboul et al., 2024). Anisotropy affects the loss wedge less strongly as $n(\rho,\lambda_{z})$ is normally not a strong function [^15] One exception to this is our NSD model at larger $\rho$ (see figure 16), where the disk-like nature of the DF leads to the majority of orbits having $l_{z}\sim l_{\rm c}$, though at such large $\rho$ ($\gtrsim 2$) these systems would contribute little to our total rates even if the DF was isotropic. of $\lambda_{z}$ near $|\lambda_{z}|\sim 0$.
It is also interesting to compare our scattering rates to previous work. We predict a substantially lower overall rate for scattering-driven TDEs, $10^{-5}$ yr-1, than commonly reported values of order $10^{-4}$ yr-1 (see for example Generozov et al. 2018). We also see that we are only in the full loss-cone regime for small $l_{\rm t}$ and intermediate $\rho$, whilst it’s normally assumed that at large $\rho$ we should always be in the full loss-cone regime. Both of these differences stem mostly from the anisotropy of our measured $n(\rho,l)$ distributions. The fact that we are not in the full loss-cone regime at larger $\rho$ also comes from our neglecting scatterings from bulge stars (which should dominate at $\rho\gtrsim 1$) though these should have little impact on the overall rate.
@@
Figure 12: The contribution to the relative rate of disruption (top panels) and cumulative relative rates (bottom panels) for disruptions by diving orbits (left panels) and scatterings (right panels) - for systems of a given $l_{\rm t}$ (in units of [pc2 Myr-1]). We also give the approximate tidal radius ($r_{\rm t}\sim l_{\rm t}^{2}(2GM_{\rm BH})^{-1}$ where $M_{\rm BH}=4\cdot 10^{6}M_{\odot}$). For the scattering calculations we show the empty loss-cone rate (solid line) and the full loss-cone rate (dotted). We take the true rate of disruptions by scattering to be whichever is smaller, and in the cumulative rate the solid or dotted regions correspond to which rate is relevant in that region. Vertical lines show the $\rho$ which contributes half of the relative rate (giving a rough idea of where we expect disrupted systems to originate from).